OpenStudy (anonymous):

Find x if y=1/2 4x/y=2(x+1)

OpenStudy (anonymous):

whats the equation?

OpenStudy (anonymous):

4x/y=2(x+1)

OpenStudy (anonymous):

y=1/2

OpenStudy (anonymous):

i believe x =2

OpenStudy (anonymous):

4x/1/2=2x+2 4x/1/2-2x=2 4x-2x=4 2x=4 x=2

OpenStudy (anonymous):

do you understand or do you want me to explain it more?

OpenStudy (anonymous):

no thank you very much